OpenStudy (anonymous):

How do you Write 4152 in its expanded form.

OpenStudy (anonymous):

4,000+100+50+2. Is that what your looking for?

OpenStudy (akashdeepdeb):

Do ti just like that other previous question. Thousands Hundreds Tens Ones 4 1 5 2 Expanded form: 4*1000 + 1*100 + 5*10 + 2*1 Understood? :)

OpenStudy (anonymous):

it has the partheness and to the second power like my past one

OpenStudy (akashdeepdeb):

1000 = 10^3 100 = 10^2 10 = 10^1 1 = 10^0 Okay?

OpenStudy (anonymous):

I hate these sorry guys

OpenStudy (anonymous):

ok